OpenStudy (anonymous):

Evaluate I= integral of (2x^2+3x+4)/(x^2+6x+10)dx pls hlp !!!!

OpenStudy (lalaly):

first do long divison

OpenStudy (anonymous):

\[\int\limits{\frac {(2x^{2}+3x+4)}{(x^2+6x+10)}}dx\]

OpenStudy (shubhamsrg):

easier should be that first take 2 common from numerator,,then brake it into denominator + (something) that something has 1 degree less than denominator.. now many methods to do this..you may use partial differentiation..

sam (.sam.):

long division you'll get \[\int\limits \left(\frac{-9 x-16}{x^2+6 x+10}+2\right) \, dx\]

OpenStudy (anonymous):

how to solve -9x-16 /x^2+6x+10

OpenStudy (anonymous):

to get \[\frac{11}{x^2+6x+10} - \frac{9(2x+6)}{2(x^2+6x+10)}\]\

OpenStudy (anonymous):

in my txt book we have a formula like \[\int \frac{(px^2+qx+r)}{(ax^2+bx+c)} =A(ax^2+bx+c)+B d/dx(ax^2+bx+c) +c \]\

OpenStudy (anonymous):

consider integral of form \[\int \frac{px^2+qx+r)}{ax^2+bx+c)}\]\ then let px^2+qx+r =A(ax^2+bx+c)+B d/dx(ax^2+bx+c) +c

OpenStudy (anonymous):

i substituted in this form but i dnt know how to find A B and C

OpenStudy (anonymous):

i.e let \[\ 2x^2+3x+4 =A(x^2+6x+10) + B d/dx(x^2+6x+10)+c\]\

OpenStudy (anonymous):

i.e \[\ 2x^2+3x+4=A(x^2+6x+10)+B (2x+6) +c\]\

OpenStudy (anonymous):

if i substitute x=(-3) we get \[\ A+C=13\]\

OpenStudy (anonymous):

what is the nxt step aftr this??? to find A ,B and C pls hlp

OpenStudy (freckles):

do you know trig sub?

OpenStudy (anonymous):

u mean the formulas

OpenStudy (anonymous):

\[\int\frac{dx}{a^2+x^2} = \frac {1}{a} tan^{-1} \frac{x}{a} +c\]\

OpenStudy (freckles):

\[\int\limits_{}^{}\frac{-9x-16}{x^2+6x+10} dx\] \[\int\limits_{}^{}\frac{-9x-16}{(x^2+6x)+10} dx\] \[\int\limits_{}^{}\frac{-9x-16}{(x^2+6x+9)+10-9} dx\] \[\int\limits_{}^{}\frac{-9x-16}{(x+3)^2+1} dx\]  \[\tan(\theta)=\frac{x+3}{1}\] \[\tan(\theta)=x+3\] differentiating both sides \[\sec^2(\theta) d \theta =1 dx\] So we have \[\int\limits_{}^{}\frac{-9(\tan(\theta)-3)-16}{\tan^2(\theta)+1} \sec^2(\theta) d \theta \] \[\int\limits_{}^{}\frac{-9(\tan(\theta)-3)-16}{\sec^2(\theta)} \sec^2(\theta) d \theta \] \[\int\limits_{}^{}[-9(\tan(\theta)-3)-16] d \theta \] did i make a mistake the rest should be easy

OpenStudy (anonymous):

no not this type

OpenStudy (anonymous):

my question s how to find the value of A B and C

OpenStudy (anonymous):

pls hlp

OpenStudy (anonymous):

sustituting x=0 we get \[\ 10A+6B+C=4\]\

OpenStudy (anonymous):

pls hlp

OpenStudy (anonymous):

sam! as u said long division method aftr long division wat we have to do

sam (.sam.):

\[\int\limits \left(\frac{-9 x-16}{x^2+6 x+10}+2\right) \, dx\] Separate \[\int\limits \frac{-9 x-16}{x^2+6 x+10} \, dx+\int\limits 2 \, dx\] Separate \[\int\limits 2 \, dx+\int\limits \left(\frac{11}{x^2+6 x+10}-\frac{9 (2 x+6)}{2 \left(x^2+6 x+10\right)}\right) \, dx\] Separatte \[11\int\limits \frac{1}{x^2+6 x+10} \, dx-\frac{9}{2}\int\limits \frac{2 x+6}{x^2+6 x+10} \, dx+\int\limits 2 \, dx\] Let \[u=x^2+6 x+10~~ and ~~ u=2 x+6\] \[11\int\limits \frac{1}{x^2+6 x+10} \, dx-\frac{9}{2}\int\limits \frac{1}{u} \, du+\int\limits 2 \, dx\] Complete the square then integration by trig substitution \[-\frac{9}{2}\int\limits\limits \frac{1}{u} \, du+11\int\limits\limits \frac{1}{(x+3)^2+1} \, dx+\int\limits\limits 2 \, dx\] We'll call "s" as another variable , \[11 \tan ^{-1}(s)-\frac{9}{2}\int\limits \frac{1}{u} \, du+\int\limits 2 \, dx\] Integrate all \[11 \tan ^{-1}(s)-\frac{9 \ln (u)}{2}+2 x+c\] Result \[-\frac{9}{2} \ln \left(x^2+6 x+10\right)+2 x+11 \tan ^{-1}(x+3)+c\]

OpenStudy (anonymous):

how did u seperate \[\frac{(-9x-16)}{(x^2+6x+10) }\]\

sam (.sam.):

\[\frac{-9 x-16}{x^2+6 x+10}\] You multiply 2 from numerator and denominator

OpenStudy (anonymous):

how ?? i didnt understand

sam (.sam.):

wait

sam (.sam.):

(-16-9x)/(x^(2)+6x+10) (11-9(x+3))/(x^(2)+6x+10) (11)/(x^(2)+6x+10)-(9(x+3))/(x^(2)+6x+10) (11)/(x^(2)+6x+10)-((9(x+3))/(x^(2)+6x+10)) (11)/(x^(2)+6x+10)-((18(x+3))/(2(x^(2)+6x+10))) (11)/(x^(2)+6x+10)-(((18)(x+3))/(2(x^(2)+6x+10))) (11)/(x^(2)+6x+10)-((9(2(x+3)))/(2(x^(2)+6x+10))) (11)/(x^(2)+6x+10)-((9(2x+6))/(2(x^(2)+6x+10)))

sam (.sam.):

@aafrin1410

sam (.sam.):

\[\frac{-16-9x}{x^{2}+6x+10}\] \[\frac{11-9(x+3)}{x^2+6x+10}\] \[\frac{11}{x^2+6x+10}-\frac{9(x+3)}{x^2+6x+10}\] \[\frac{11}{x^2+6 x+10}-\frac{9 (2 x+6)}{2 \left(x^2+6 x+10\right)}\]

OpenStudy (anonymous):

thank you